# DynamicPAE: Generating Scene-Aware Physical Adversarial Examples in Real-Time

### v1

@@ 4.3 On Overcoming the Noisy Gradient in PAE Training
Existing works have shown the special characteristics of AEs and the adversarial attack problems, including its physical robustness [18], distributional properties [60, 26, 61], and gradient properties of the optimization process [25]. For the dynamic PAE training, we identify the key challenge as the noisy gradient problem, causing the problem of degenerated solution and latent evasion in the generative training, which can be regarded as originating from the necessary randomness injected into PAE training. To illustrate the challenge and evaluate the effectiveness of the proposed model and solution, we analyze each problem and our corresponding solution in terms of gradients and sample/latent space.
4.3.1 Analysis and Solving the Noisy Gradient Problem
Obfuscated Gradient Perspective: Obfuscated Gradient [25] is recognized as a typical defense mechanism in digital adversarial attacks since the defenders could significantly increase the difficulty of performing gradient descent optimization by adding random transformation. The solution for the attack is applying multiple-step attacks, or EoT, which is also a key algorithm for PAE [29] since the fluctuation in the physical world can be regarded as naturally introduced randomness.
First, we show that the simulated physical adversarial patch attacks, as a typical physical attack, can be regarded as facing the same obfuscated gradient, or noisy gradient problem indicated by our experimental results, as the spatial transformations significantly increase the difficulty of PGD attacks. Specifically, as shown in Table V and Table VI, an increase in the total number of steps, whether it’s the number of EoT iterations or the total iteration steps of the optimizer, consistently improves the PGD performance under spatial transformations. Additionally, we use 2048 steps in the main experiment based on these results.
TABLE V: Comparison with PGD in the patch attack on Yolo-v5s, evaluated with COCO dataset.
Method EoT steps Optim. steps $AP_{50}$ $AP_{01}$ time
Clean 79.0% 87.6%
PGD 1 128 50.4% 65.7% 18.6s
4 128 38.1% 55.2% 73.8s
1 512 38.0% 54.5% 74.1s
16 128 29.0% 45.2% 230.3s
4 512 27.5% 43.3% 295.2s
1 2048 27.0% 42.3% 296.5s
DynamicPAE 16.3% 28.8% 0.1s
TABLE VI: Comparison with PGD in the patch attack on Yolo-v8n, evaluated with COCO dataset.
Method EoT steps Optim. steps $AP_{50}$ $AP_{01}$ time
Clean 86.4% 92.1%
PGD 1 128 73.8% 83.5% 11.4s
4 128 65.9% 77.6% 44.9s
1 512 65.1% 77.1% 45.5s
16 128 60.2% 73.9% 140.1s
4 512 58.4% 71.7% 180.0s
1 2048 57.7% 71.5% 181.5s
DynamicPAE 27.9% 37.8% 0.1s
The noisy gradient problem also brings severe challenges for training the dynamic generator, and our methodology focused on it. We present the performance of DynamicPAE framework after solving the problem. To ensure a fair comparison, both our model and PGD are evaluated by resampling the patch locations. Note that even with this high step count, the performance of PGD is still less aggressive than our result. In the main performance evaluation, we found that the Zero transformation settings facilitate the PGD method in achieving better results. We further analyze it on the Inria dataset. As shown in Table VII, PGD surpasses our method in attack capability at 32 steps, but our method still retains an advantage in inference time. From another perspective, the superiority of our method mainly lies in overcoming physical-world transformations.
@@ 4.3.1 Analysis and Solving the Noisy Gradient Problem
TABLE VII: Comparison with PGD in the patch attack on Yolo-v5s, evaluated with zero transformation.
@@
PAE Sample Space Perspective: When training PAE generators, we find that the main challenge caused by the noisy gradient is the degenerated solution presented as mode collapse, and the theoretical modeling is constructed in section 3.1. We visualized the training result of the generated PAEs on the setting of Yolov5, base transformation, and COCO dataset. The results of training without the auxiliary residual task are shown in Figure 6a. Although these samples achieve attack capabilities, they exhibit significant mode collapse with only pixel-level weak differences between different samples. Figure 6b illustrates the results under the same configuration after introducing the residual task. The first row, the second row, and the last two rows represent the case with attack weight $\lambda=0$, $\lambda=1$, and $\lambda$ sampled between $[0,1]$, respectively. It can be observed that the diversity of the generations is significantly enhanced.
By connecting with the proposed model, the experiment shows that: ❶ the formulated limited feedback information restriction, which models the noisy gradients and predicts the degeneracy problem of the conditional PAEs, matches the experiment results. ❷ By breaking the restriction of the limited feedback information restriction, the residual-task guided training overcomes the degeneration problem, which simultaneously demonstrates the effectiveness of the noise gradient model and method.
@@
Figure 6: Comparison of patch generation.
@@
In addition, this mode collapse is different from the posterior collapse in variational autoencoders (VAE), in which the noise is injected in latent $\mathbf{Z}$ for prior-based regularization instead of stemming from the task. Moreover, although diversity is not the first goal of the attack tasks, this mode collapse is weakening the attack performance to be close to static PAEs, thus we recognize it as the degeneration problem. The comparison of the attack performance is shown in ablation studies.
4.3.2 Analysis and Solving the Latent Evasion Problem
Gradient Explosion Perspective: Another difference from training generative model on benign data is that the learning of conditional PAE autoencoder leads to infinite gradient, and we recognize it as latent evasion. Experiments are conducted on identical model structures with different tasks, and the $l_{\infty}$-norm of parameter gradients are measured. The result is shown in Figure 7. Denoted as Joint, the maximum gradient begins to grow uncapped after 30K steps until numeric overflow. However, denoted as Clean and Attack, the $l_{\infty}$-norm curves of gradients remain normal when only fitting the PAEs and the clean sample. Thus the infinite gradient is a result specific to fitting diversed PAEs and clean data. We further analyze the maximum gradient and found that the gradient is located at the patch decoding network $Dec:\mathcal{Z}\to\mathbb{D}$. We argue that even though the image is generated by constraining the output to the interval $[0,1]$, the hidden layer features remain unconstrained during modeling diverse PAEs, leading to the emergence of the latent evasion. Related theoretical analysis also shows the possibility of the infinite gradient when minimizing the VAE energy function [62]. We heuristically regularize the latent representation $\boldsymbol{z}_{adv}$. As shown in the curve Regularized Joint in Figure 7, it solves the problem of infinite gradient.
Figure 7: Inf. norm of parameter gradients in different tasks. In the case of Joint the gradient exploded while the regularization solves the problem.
Latent Space Perspective: We provide further visualized analysis on the latent space to illustrate the latent evasion problem and the regularization in the supplementary material. To interpret the learned correlation, we perform analysis based on the latent representation $\mathbf{Z}=\mathrm{Enc}(\mathbf{P}_{\mathrm{X}})$. The visualized results are shown in the supplementary material, indicating that the model successfully learns the representation of PAEs and their correlation with the observed scenes.
4.4 On the Practicality in Agnostic Attack Scenarios
@@
To evaluate the real-world practicality of DynamicPAE when facing agnostic scenarios, we conduct experiments on different environments and parameters to show the effectiveness of the proposed simulation model. The experiments demonstrate the ability of DynamicPAE to (1) reveal potential security issues and (2) generate adversarial evaluation data that is aligned with the real world in applications.
Results in Physical Environments: We implement our model and deploy it with edge-computing devices (2080Ti) and the local area network (LAN), and evaluate the attack performance of the framework within the physical environment. As shown in Figure 8, our method generally achieved higher attack performance. Although our attacks are experimental, real attacks can be realized with dynamic materials (e.g., color-changing costumes [63]) and spread through transformation simulation. Furthermore, we find that presenting the static patch in varying conditions could also the attack effectiveness, indicating that the adversarial examples generated by our method achieve a balance between attack adaptability and robustness, and the visualization, including more cases showing the adaptability and the robustness, is provided in the supplementary material.
@@
Confirmation of Physical Robustness:   Since accurate location as model input may be difficult to obtain in applications, we perform tests with inaccurate inputs. As shown in Table VIII, the attack performance did not drop much as we resample the patch locations, which is attributed to the implicit regularization of DNN and augmentation pipeline. Moreover, the attack performance is still significantly higher than the baselines in the main performance evaluation. This indicates that the model can leverage precise location information when available, while still maintaining relatively high performance even with inaccurate location inputs.
TABLE VIII: Evaluation of the robustness against imprecise location inputs targeting Yolo-v5s on the COCO dataset. Randomized refers to re-sampling the attack locations.
@@
Results in Simulated Environment: To evaluate the applicability of the proposed model in the simulated testing, we obtain the simulation results from different perspectives in the CARLA autonomous driving simulation platform [64], divide them into training and testing datasets, and apply the adversarial example as a perturbation of the vehicle’s texture. Supplement to the digital environment benchmark, we compare the performance with the GAN-based perturbation generation framework AdvGAN, since both our model and AdvGAN are learning-based and real-time.
As shown in Table IX, our model also achieves strong perturbation attack performance and significantly enhances the attack aggressiveness compared to AdvGAN. To measure the stealthiness of the attack more accurately, we measure the SSIM similarity in addition to the $l_{\infty}$ constraint. It can be seen that the attack performance of our method significantly outperforms the baseline method for all different magnitude settings. Interestingly, our method performs better in the transformed view than in the identical view when $\varepsilon=1\%$, which may be attributed to the increased generalization due to the increased task complexity, suggesting that our method is more suitable for complex physical world attacks.
TABLE IX: Evaluation results on CARLA simulation. The illustration of different view settings is shown in Figure 9.
Identical View Transformed View
$\varepsilon$ Method $AP_{50}$ $AP_{01}$ SSIM $AP_{50}$ $AP_{01}$ SSIM
0% Clean 43.23% 46.56% 1.0000 51.08% 52.56% 1.0000
1% AdvGAN 29.67% 38.40% 0.9983 34.30% 44.17% 0.9994
1% DynamicPAE 5.35% 12.39% 0.9989 3.31% 7.53% 0.9994
5% AdvGAN 6.0e-5 7.0e-3 0.9888 2.9e-3 1.1e-2 0.9928
5% DynamicPAE $<$1e-6 2.5e-6 0.9969 $<$1e-6 8.5e-6 0.9931
@@
Hyper-parameter Sensitivity: Many generative models are sensitive to hyperparameters, especially GAN models, which harms their practicality. To make it clear, we present an investigation of the hyperparameters in our proposed model. The analysis includes the parameter $\alpha_{0}$ as the initial weight of invisibility and $\gamma$ in the regularization loss in Eq. 11. The results are shown in Table X. For regularization loss, we evaluate its weight with settings $\gamma\in\{10^{-4},10^{-2},1.0\}$, and the attack performance remains steady, indicating that it does not require excessive parameter tuning. For the task scheduler, we evaluate it with different initial weights of invisibility $\alpha_{0}$, representing the variability in loss intensity between attack tasks, and both the auto-adjusted weight $\alpha_{final}$ and evaluation results remain similar, indicating the functionality of the skewness-based objective re-weighting mechanism that eliminates the free parameter and strike a consistent balance between attack and stealthiness. Moreover, on the aggressiveness-distortion curves of the main performance experiments, the uniform and consistent data point distribution of our method across the different models also indicates it.
@@
TABLE X: Analysis of parameter sensitivity. The weight of regularization $\gamma$ is insensitive and the sensitivity of the weight of the residual task is reduced by the skewness-aligned objective re-weighting method.
@@ 4.5 Ablation Studies
Auxiliary Residual Task: Ablation experiments across different target models and different patch transformations are conducted to evaluate the residual task. $\lambda$ is set to 1 during tests. As shown in Figure XI, the constructed auxiliary residual task consistently improves the attack performance of dynamic PAE, demonstrating the necessity of such construction in the dynamic PAE optimization process. We recognize that the improvement is more significant in large victim models that are hard to attack, including Yolo-v5m and Faster-RCNN, which may be because larger models have more diverse vulnerabilities, allowing for the discovery of more conditional solutions through the improved optimization process. We also recognize that in the setting of P+, where the spatial transformation in attack injection $\oplus$ is more diverse, the drop ratio of $AP$ is relatively low. This may be because of higher uncertainty makes the information available for adaptation limited, with less room for performance improvement. It is worth exploring how to construct more effective residual tasks or alter attack tasks to further enhance attack capabilities.
TABLE XI: Ablation study of the auxiliary residual task. By collaborating with the attack training, the residual task consistently improves the attack performance compared to vanilla training (in columns w/o).
@@
Latent Regularization: We compare the differences among regularization in terms of the attack performance. As shown in Table XII, the center-regularized $\boldsymbol{z}_{adv}$ achieves the optimal performance among different settings and solves the problem of gradient explosion when learning diverse PAEs through the auxiliary residual task. Note that without the auxiliary residual task (Aux. Task), the PAE decoder $\mathrm{Dec}$ degenerates to a single static PAE. We noticed classical VAE regularization still suffers the infinite gradient in the model of Yolo-v8n and the attack performance on Yolo-v5 is lower, indicating the widely-used Gaussian prior might not be a proper selection for the latent of explored PAEs.
TABLE XII: Ablation of latent regularization. Encouraging the codewords of PAEs to be located in the center of the representation space achieves better attack performance and training stability.
@@
$l_{2}$-norm (DynamicPAE Applied) v5 29.34% 48.59%

### v3

@@ 4.3 On Overcoming the Noisy Feedback in Training
Existing works have shown the special characteristics of AEs and the adversarial attack problems, including AE’s physical robustness [22], distributional properties [56, 19, 57], and gradient properties of the optimization process [18]. For the dynamic PAE training, we identify the key challenge as noisy feedback, which causes the problem of degenerated solution. We model the problem with the limited feedback information restriction and construct the technical solution based on it. This section explores the impact of noisy gradient feedback and the effectiveness of our model.
$\bullet$ Noisy Feedback Causes PGD Attacks Time-Consuming: We first show that the simulated physical adversarial patch attacks, as a typical physical attack, can be regarded as facing the obfuscated gradient [18], or noisy gradient feedback problem indicated by our experimental results, as the spatial transformations significantly increase the difficulty of PGD attacks. Obfuscated gradient has been recognized as a typical adversarial defense mechanism that increases the difficulty of performing optimization by adding a random transformation. The randomness in the physical simulation, introduced as the EoT [23] transformation in PAE methods, has the same effect. Specifically, when randomness is not injected, Table VIII shows that PGD achieves high attack performance when the step number reaches 32. However, when randomness is injected with Base transformation in Table IX, an extensive number of total steps, whether it’s EoT iteration steps or the optimization steps, is needed to achieve satisfactory attack performance. This illustrates the challenges of PGD optimization in the presence of noisy gradient feedback resulting from random simulation.
@@
TABLE VIII: Comparison with PGD in the patch attack on Yolo-v5s, evaluated with zero transformation in Table III and Inria dataset.
@@
TABLE IX: Comparison with PGD in the patch attack on Yolo-v5s and Yolo-v8n, evaluated with base transformation in Table III and more complex COCO dataset.
EoT
Steps
Optim.
Steps
Yolo-v5s Yolo-v8n
Method $AP_{50}$ $AP_{01}$ time $AP_{50}$ $AP_{01}$ time
Clean 79.0% 87.6% 86.4% 92.1%
PGD 1 128 50.4% 65.7% 18.6s 73.8% 83.5% 11.4s
4 128 38.1% 55.2% 73.8s 65.9% 77.6% 44.9s
1 512 38.0% 54.5% 74.1s 65.1% 77.1% 45.5s
16 128 29.0% 45.2% 230.3s 60.2% 73.9% 140.1s
4 512 27.5% 43.3% 295.2s 58.4% 71.7% 180.0s
1 2048 27.0% 42.3% 296.5s 57.7% 71.5% 181.5s
T-sea 48.2% 63.1% 60.4% 75.0%
DynamicPAE 16.3% 28.8% 0.1s 27.9% 37.8% 0.1s
We present the performance of static PAE (T-sea) and DynamicPAE framework for comparison. To ensure a fair comparison, both our model and PGD are evaluated by resampling the patch locations (same setting as Table XI). The result shows that PGD shall take hundreds of steps to surpass the performance of static PAE under the base transformation, indicating its incapability of real-time attacks.
$\bullet$ Noisy Feedback Causes Generative Training Degenerate: When training PAE generators, we find that the new challenge caused by the noisy gradient feedback is the degenerated solution presented as mode collapse, and the theoretical modeling is constructed in section 3.1. We visualized the training result of the generated PAEs on the setting of Yolov5, base transformation, and COCO dataset. The results of training without the auxiliary residual task are shown in Figure 9a. Although these PAEs achieve attack capabilities, they exhibit significant mode collapse with only pixel-level weak differences between different PAEs. Figure 9b illustrates the results under the same configuration after introducing the residual task. The first row, the second row, and the last two rows represent the case with attack weight $\lambda=0$, $1$, and sampled between $[0,1]$, respectively. It can be observed that the diversity of the generations is significantly enhanced, especially when $\lambda=1$ (row 2 of Figure 9a and 9b), where the training objectives are equivalent.
@@
Figure 9: Comparison of Patch Generation. After applying the residual task, the mode collapse problem in training the dynamic PAE generator is solved.
Figure 10: Patch Visualization for Different Targets. The original training results in identical patches, while the residual-guided training results in different patches.
We visualize the patch generation results for other target models in Figure 10. Without the residual-guided training, the generated patches are identical in all four target models. Residual-guided training successfully finds diverse solutions when attacking each model. Although the explored PAE patterns are not diverse, the escape behavior of the degenerated and identical solution is consistent. Exploring more diverse aggressive solutions might be a difficult task within limited training computation, as finding each pattern requires a large number of queries under the assumption of limited feedback information.
@@
In addition, this mode collapse is different from the posterior collapse in variational autoencoders (VAE), in which the noise is injected in latent $\mathbf{Z}$ for prior-based regularization instead of stemming from the task. Moreover, although diversity is not the primary goal of the attack tasks, this mode collapse is weakening the attack performance of dynamic PAEs to be close to static PAEs, thus we recognize it as the degeneracy problem.
$\bullet$ Residual-Guided Training Overcomes the Degeneracy Problem: To show the different PAEs in Figure 9b and 10 is generated based on the attacker’s observation rather than randomly generated, we further analyze the latent representation $\mathbf{Z}$ of the trained generator. We first annotate the latent representations based on the style of PAEs, and apply LDA for dimension reduction. The result is shown in Figure 11a, indicating that the model successfully learns a linearly decomposable representation of PAEs. We conduct the K-nearest neighbor (KNN) search (k=3) in the dimension-reduced space and visualize the corresponding physical context in Figure 11b. The outcomes demonstrate that the KNN result has significant similarity with the KNN query in terms of certain characteristics, including human actions, the chromatic attributes of clothing, and the exposed body parts et al., indicating that our model successfully learns a distinction between $\mathbf{P}_{\mathrm{X}}$ based on the vulnerability of the target model by learning PAE generation.
(a) PAE Representation Space
(b) Physical Contexts KNN
Figure 11: Visualization of the learned correlation between PAEs and the related physical environments (collected by latent-space K-nearest neighbor (KNN)).
The comparison of the attack performance is shown in ablation studies. In addition to object detection tasks, [55] attempted the generative approach for real-time digital adversarial patch attacks against image classification and face recognition. We observe that, even without incorporating randomized physical simulation and with patch sizes limited to $\times32$, it still suffers from certain degeneracy issues. We introduce the proposed training strategy into this model (detailed implementation is in supplementary material) and test it with the 2%-ratio patch attack setting and the VGGFace classification task. As shown in Figure 12 and Table 12, the final ASR and training convergence are significantly improved ($10\%$ ASR, $4\times$ training speed). This highlights the criticality of the proposed training technique.
Overall, the experimental analysis shows that: ❶ The obfuscated gradient problem exists in the PAE training environment. ❷ limited feedback information restriction, which models the noisy gradients and predicts the degeneracy problem of the conditional PAEs, matches the experiment results. ❸ By breaking the restriction of the limited feedback information, the residual-task guided training overcomes the degeneration problem. This result simultaneously demonstrates the effectiveness of the limited feedback information restriction model and the residual-task construction.
Figure 12: Loss curves under different training strategies.
TABLE X: ASR on the VGGFace validation set. The residual task improves the training convergence.
Epoch GDPA + Res. Task
1 5.1% 58.9%
2 6.2% 70.8%
4 20.9% 77.8%
25 57.1% 88.2%
50 42.3% 88.0%
100 78.9% 88.6%
4.4 On the Practicality in Real Attack Scenarios
@@
To evaluate the real-world attack capability of DynamicPAE, we conduct experiments on different environments and parameters to show the effectiveness of the proposed alignment method. The experiments demonstrate the ability of DynamicPAE to (1) reveal potential security issues and (2) generate PAEs that are robust in the more complex deployment environments.
$\bullet$ Results in Physical Environments: We implement our model and deploy it on laptop, and evaluate the attack performance of the framework within the physical environment. As shown in Figure 13, our method achieves higher attack performance since neither of the two static PAEs caused the target detector to miss the detection. Additional physical experiments, accompanied by visualized figures and videos that demonstrate real-time adaptability to lightning changes and robustness against physical disturbances, are provided in the supplementary material.
@@
$\bullet$ Confirmation of Physical Robustness: Since an accurate location, as the input of $\mathcal{G}$, may be difficult to obtain in applications, we perform tests with inaccurate inputs. During the training, the patch location is encoded into the generator’s input $\mathbf{P}_{X}$, and patches’ positioning parameters $\boldsymbol{\theta}$ are drawn from the distribution $p$ defined by Table III. In Table XI, Randomized denotes the patch locations corresponding to $\boldsymbol{\theta}$ and $\mathbf{P}_{X}$ are i.i.d. parameters drawn from $p$, or the patch position and the generator’s positioning input are unrelated, and Accurate Input denotes they share identical positioning parameters for each person attack test case. As the table shows, the attack performance did not drop much as we resampled the patch locations, which is attributed to the implicit regularization of DNN and the augmentation pipeline. Moreover, the attack performance is still significantly higher than the baselines in the main performance evaluation. This indicates that the model can leverage precise location information when available while still maintaining relatively high performance even with inaccurate location inputs.
TABLE XI: Evaluation of the robustness against imprecise location inputs targeting Yolo-v5s on the COCO dataset. Randomized refers to re-sampling the attack locations.
@@
$\bullet$ Generalization in PAE Deployments: Unlike the generalization topic centered on the static data in classical tasks, the generalization of PAEs shall consider the manufacturing and resampling distortions [58]. To evaluate the generalizability, we test our trained generator in the 3D simulated physical deployment scenarios [59, 60]. As shown in Figure 10, compared to our 2D training environment (e.g., AF), the patch transformations in the 3D simulation environment exhibit larger variations and are also more closely aligned with the physical world. We evaluate the trained DynamicPAE with T-sea, GAN-NAP, and AdvReal in both environments. The background dataset for training is COCO, Inria, Inria, and nuScences, respectively. nuScences [61] is a more realistic dataset for autonomous driving. For the physical observation, only a coarse patch positioning in the 3D evaluation is provided as the input of DynamicPAE.
Table XII shows the key settings and results. Besides achieving better attack performance in the 2D scenario, our DynamicPAE maintains the superiority over T-sea and GAN-NAP in the 3D scenario. Therefore, by training with the Conditional-Uncertainty-Aligned Data Model defined in section 3.1, DynamicPAE can generalize to unseen physical deployment processes. Although AdvReal achieved better results in the 3D scenario by training the static PAE with consistent 3D simulation, its performance on our small 2D-patch attack scenario does not improve. Furthermore, the improved attack performance with better simulation (Base $\to$ AF) indicates the direction of scaling DynamicPAE with higher-fidelity physical world simulation systems.
TABLE XII: Performance Evaluation on Different PAE Deployment Simulations. 2D-Base/AF denotes the settings in Table III, and 3D-AdvReal denotes the AdvReal 3D simulation setting illustrated in Figure 14.
Attack
Method
Training
Settings
ASR-Inria
2D-Base
ASR-nuScenes
3D-AdvReal
GAN-NAP 2D-Base 29.51% 14.94%
T-sea 2D (T-sea) 30.34% 64.07%
AdvReal 2D+3D (AdvReal) 28.73% 98.27%
DynamicPAE 2D-Base 55.30% 69.26%
DynamicPAE 2D-AF 67.35% 93.29%
Figure 14: PAE deployment simulations. (Left: AF defined in Table III. Right: 3D simulation in AdvReal [60]).
@@
$\bullet$ Hyper-parameter Sensitivity: The training of the generator shall be aligned with the stability requirement of the adversarial evaluation scenario. However, many generative models are sensitive to hyperparameters, especially GAN models, which harms their practicality. To make it clear, we present an investigation of the hyperparameters in our proposed model. The analysis includes the parameter $\alpha_{0}$ as the initial weight of invisibility and $\gamma$ in the regularization loss in Eq. 11. The results are shown in Table XIII. For regularization loss, we evaluate its weight with settings $\gamma\in\{10^{-4},10^{-2},1.0\}$, and the attack performance remains steady, indicating that it does not require excessive parameter tuning. For the task scheduler, we evaluate it with different initial weights of invisibility $\alpha_{0}$, representing the variability in loss intensity between attack tasks, and both the auto-adjusted weight $\alpha_{final}$ and evaluation results remain similar, indicating the functionality of the skewness-based objective re-weighting mechanism that eliminates the free parameter and strike a consistent balance between attack and stealthiness. Moreover, on the aggressiveness-distortion curves of the main performance experiments, the uniform and consistent data point distribution of our method across the different models also indicates it.
@@
TABLE XIII: Analysis of parameter sensitivity. The weight of regularization $\gamma$ is insensitive. The sensitivity of the weight of the residual task is reduced by the skewness-aligned objective re-weighting method.
@@ 4.5 Ablation Studies
TABLE XIV: Ablation study of the auxiliary residual task. By collaborating with the attack training, the residual task consistently improves the attack performance compared to vanilla training (in columns w/o).
@@
TABLE XV: Ablation study of the residual training. The improvement attributed to the proposed residual task is more significant than the residual neural architecture.
Neural Arch. Res. Task Model $AP_{50}$ $AP_{01}$
w/o All Skip
Connections
in $\mathrm{Dec}$
v5 18.57% 35.14%
v8 49.36% 54.63%
v5 17.11% 32.96%
v8 31.32% 44.26%
w/o Skip-$\mathbf{Z}$
Connections
in $\mathrm{Dec}$
v5 22.00% 40.50%
v8 44.04% 49.44%
v5 17.66% 34.22%
v8 27.11% 36.19%
All Skip
Connections
Added
v5 22.80% 41.01%
v8 40.82% 46.12%
v5 14.87% 26.72%
v8 26.04% 35.85%
$\bullet$ Auxiliary Residual Task: Ablation experiments across different target models and different patch transformations are conducted to evaluate the residual task. $\lambda$ is set to 1 during tests. As shown in Figure XIV, the constructed auxiliary residual task consistently improves the attack performance of dynamic PAE, demonstrating the necessity of such construction in the dynamic PAE optimization process. We recognize that the improvement is more significant in large victim models that are hard to attack, including Yolo-v5m and Faster-RCNN, which may be because larger models have more diverse vulnerabilities, allowing for the discovery of more conditional solutions through the improved optimization process. We also recognize that in the setting of P+, where the spatial transformation in attack injection $\oplus$ is more diverse, the drop ratio of $AP$ is relatively low. This may be because the higher uncertainty makes the information available for adaptation limited, with less room for performance improvement. It is worth exploring how to construct more effective residual tasks or alter attack tasks to enhance attack capabilities further.
Introducing residual connections in neural networks is a classic method in deep learning to address optimization issues. We further conduct experiments to compare residual connections and the residual task. The DynamicPAE framework adopts the image decoder architecture of BigGAN [49], which combines the additional Z-skip residual connections with the ResNet architecture. We perform the ablation study on Z-skip and other skip connections in ResNet, along with the proposed residual task, in dynamic PAE generation. As shown in Table XV, the residual task significantly improves performance across different configurations. The improvement from residual connections is not significant, and negative effects exist in several settings. Therefore, the residual task technique is more important than the residual connection in training the dynamic PAE generator.
TABLE XVI: Ablation of latent regularization. Encouraging the codewords of PAEs to be located in the center of the representation space achieves better attack performance and training stability.
@@
$l_{2}$-norm
(DynamicPAE
Applied)
v5 22.80% 41.01%